The difference between the square of a positive number and five times the number is 36. Find the number.
Answer by ankor@dixie-net.com(12692)

The difference between the square of a positive number and five times the number is 36. Find the number.
;
Just write an equation for what it says:
:
x^2 - 5x = 36
:
x^2 - 5x - 36 = 0; a quadratic equation
Factor to:
(x-9)(x+4) =0
It specifies a positive number, therefore:
x = +9
:
:
See if that is true:
9^2 - 5(9) =
81 - 45 = 36